OpenStudy (anonymous):

the sum of a number and 1 over 8 of the number is 54 find the number

OpenStudy (anonymous):

x+1/8x=54 combine the x(which is 9/8=make common denominators) and then move it to the other side. x=48.

OpenStudy (anonymous):

(x + 1)/8x = 54 x + 1 = 432x Can you do it like this?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

it doesnt give u the right answer...?

OpenStudy (anonymous):

Which is the right equation? (x + 1)/8 = 54 x + 1/8 = 54

OpenStudy (anonymous):